OpenStudy (anonymous):

Find the 50th term of the sequence 5, –2, –9, –16, ... –352 –343 –338 –331

OpenStudy (anonymous):

5. Find the 50th term of the sequence 5, –2, –9, –16, ... (1 point) (0 pts) –352 (0 pts) –343 (1 pt) –338 (0 pts) –331 1 /1 point

OpenStudy (campbell_st):

ok... the sequence is arithmetic so use the formula for a term in an arthmetic sequence \[T_{n} = a + (n - 1) d\] a= 1st term, n = number of terms d = common difference so \[T_{50} = 5 + (50 -1) \times (-7)\] just evaluate for the answer

OpenStudy (anonymous):

a1 would be the first term(5) nth term = 50 an=a1+(n-1)(d) a50=5+(50-1)(-7) a50=5-7(49)= -338